# Data Processing for the OpenGPT-X Model Family

### v1

@@ 5.2 Web Data
Figure 7: Percentage of data removed after deduplication for each language, showcasing the varying impact of deduplication.
Our deduplication methodology is based on a MinHash + Locality Sensitive Hashing approach that is explained in detail in [34], where the choice of algorithm, along with precision and recall metrics, are analyzed. In this section, we shift focus to examining how deduplication disproportionately affects different languages. Figure 7 illustrates the percentage of data removed after deduplication for each language, ranging from a minimum of 29.03% for Hungarian to a maximum of 68.95% for English. While it may seem intuitive to correlate the deduplication percentage with data availability, a Pearson correlation analysis reveals only a modest positive correlation of 0.54 (p = 0.006).
This suggests that some languages may be subject to an “unfair” amount of deduplication relative to their size. Although deduplication operates by identifying and removing similar content, this implies that certain languages contain a higher proportion of similar data, which may result in a final model that struggles to capture linguistic nuances.
To evaluate this disproportionality, we introduce the Deduplication Disparity Index (DDI), which measures the impact of deduplication on each language relative to its available web data. For each language $l$, the DDI is calculated as the Z-score of the deduplication-to-data ratio $R_{l}$, which measures the percentage of data deduplicated in relation to the total web data available for that language:
$\operatorname{DDI}_{l}=\frac{R_{l}-\mu_{R}}{\sigma_{R}},\quad\text{where}\quad R_{l}=\frac{d_{l}}{W_{l}}$
Here, $d_{l}$ is the deduplication percentage for language $l$, $W_{l}$ is the total number of web data words for language $l$, $\mu_{R}$ is the average deduplication-to-data ratio across all languages, and $\sigma_{R}$ is the standard deviation of the deduplication-to-data ratios. This Z-score transformation centers the $\operatorname{DDI}$ around zero, highlighting languages where deduplication has a disproportionately high or low impact.
Languages with a high positive Z-score, such as Maltese ($\operatorname{DDI}_{mt}=4.519$) and Croatian ($\operatorname{DDI}_{mt}=1.0735$), are significantly impacted by deduplication, while others have scores closer to zero ($\operatorname{DDI}=-0.25\pm 0.012)$, indicating a more balanced deduplication-to-data ratio. This metric provides a systematic way to assess fairness in the data processing pipeline and highlights languages that might require further attention during preprocessing.
@@ 5.2.1 Summary
Deduplication was the most resource-intensive step, consuming 80.8% of total CPU hours allocated for data processing. We introduced the Deduplication Disparity Index (DDI) to identify languages disproportionately affected by deduplication, highlighting the need for tailored processing strategies to ensure fairness and data quality.
@@ 8.3.2 List of Curated Datasets
ampsmath_khan EN JSON / TXT MIT Math 102.985 12.081 0%
ampsmath_mathematica EN JSON / TXT MIT Math 4.824.189 232.502 0%
cdrs_bt DE TXT CC0 1.0 Law & Admin. 124.487 515.053 8%
ce_bag DE TXT CC0 1.0 Law & Admin. 5.290 18.487 0%
ce_bfh DE TXT CC0 1.0 Law & Admin. 10.310 22.635 0%
ce_bgh DE TXT CC0 1.0 Law & Admin. 71.976 104.002 0%
ce_bpatg DE TXT CC0 1.0 Law & Admin. 30.772 66.029 0%
ce_bverfg DE TXT CC0 1.0 Law & Admin. 8.449 20.679 0%
ce_bverwg DE TXT CC0 1.0 Law & Admin. 27.099 49.664 0%
cpp_bt DE TXT CC0 1.0 Law & Admin. 4.087 235.502 0%
@@
cd_icj EN, FR TXT CC0 1.0 Law & Admin. 4.565 25.979 0%
cd_pcij EN TXT CC0 1.0 Law & Admin. 518 2.117 0%
dagw DA TXT CC BY 4.0 Web 285.634 922.700 5%
@@
dcep EU24 TXT EUPL Law & Admin. 1.227.996 1.117 5%
medi_notice DE, FR, IT TSV CC BY SA Medical 22.219 52.068 0%
enc2021 EU24 VERT CC BY-NC 4.0 Web 4.361.539 1.004 5%
@@
emea EU24 TXT CC BY 4.0 Medical 17.960 152.252 4%
europarl EU24 XML CC0 1.0 Law & Admin. 63.937 660.716 0%
eurlex EU24 JSONL CC BY 4.0 Law & Admin. 4.463.480 13.235 8%
ecb_corpus EU24 XML CC0 Finance 19 53.894 0%
@@
german_legal_cases DE JSONL CC0 1.0 Law & Admin. 249.240 749.060 0%
german_political_speeches DE XML CC BY-SA 4.0 Law & Admin. 6.659 11.366 0%
greek_legal_code EL JSONL CC BY 4.0 Law & Admin. 40.929 28.791 1%
@@
korpus_malti MA JSONL CC BY-NC-SA Web 103.874 276.204 11% [^29] Noisy data and short document.
@@
marcell BG, HR, HU, RO, SL, SK, PL CONLLUP / TXT CC0 1.0 Law & Admin. 280.629 373.283 5%
macocu EU24 XML CC0 1.0 Web 79.164.225 27.930 12% [^30] Short documents.
multi_un EU24 XML CC BY-NC Law & Admin. 262.391 1.096 0%
multi_eurlex EU24 JSONL CC BY-NC-SA 4.0 Law & Admin. 1.038.305 1.184 1%
medical_t5 DE, EN, ES, FR JSON Apache-2.0 Medical 94.269 90.741 1%
ncc NO JSON Various [^31] The ncc corpus includes the following licenses : NLOD 2.0, CC0 1.0, CC BY-NC 2.0, CC BY-SA 3.0 Knowledge Base 1.241.694 2.920 60% [^32] Short documents.
openlegaldata DE JSONL CC0 1.0 Law & Admin. 103.870 363.962 0%
opensubtitles2018 EU24 XML CC BY-NC Culture 2.257.513 1.398 22% [^33] Noisy documents.
open_web_math EN Parquet ODC-By 1.0 Math 5.956.611 7.378 2%
paisa IT TXT / XML CC BY-NC-SA 3.0 Web 318.730 207.382 7%
parlamint EU24 XML / TEI CC BY Law & Admin. 40.995 1.095 3%
pes2o EN JSONL ODC-By 1.0 Academic 38.958.175 42.172 0%
pile_v2_freelawopinions EN JSONL CC BY-ND 4.0 Law & Admin. 4.410.012 10.408 0%
pile_hackernews EN JSONL CC BY-NC Web 331.205 246.149 1%
pile_nih_exporter EN JSONL CC0 1.0 Medical 930.048 307.055 0%
pile_openwebtext2 EN JSONL MIT Books 15.752.217 10.633 1%
pile_v2_philarchive EN JSONL Various [^34] This Pile part is extracted from ArXiv where the author can decide between a variety of licenses. Academic 40.344 487.032 1%
pile_pmc_abstracts EN JSONL Various [^35] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 15.476.085 3.173 0%
pile_pmc_extracts EN JSONL Various [^36] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 2.808.849 12.113 1%
ppc PL XML CC BY 4.0 Law & Admin. 38.304 472.193 1%
@@
rp_arxiv EN JSONL Apache-2.0 Academic 1.503.469 10.202 3%
setimes_corpus BG, BS, EL, EN, HR, MK, RO, SQ TXT CC BY-SA 3.0 News 246.206 69.279 7%
seimas EN, LT XML / TMX CC BY 4.0 Law & Admin. 3540 47.319 0%
@@
slwac SL CONLLU CC BY-SA 4.0 Web 1.484.546 953.803 16% [^38] Low language score due to similar languages.
spanish_legal_corpora ES TXT CC BY 4.0 Law & Admin. 15 1.383.749 0%
@@
starcoder_data XX JSONL Various [^40] The Stack is a collection of source code from repositories with various licenses. Any use of all or part of the code gathered in The Stack must abide by the terms of the original licenses, including attribution clauses when relevant. Source Code 206.634.734 73.064 0%
swiss_judgment_prediction DE, FR JSONL CC BY-SA 4.0 Law & Admin. 262.789 144.599 0%
swiss_legislation_corpus DE, FR TXT CC BY-SA Law & Admin. 2 10.407 0%
swiss_policy_documents DE, FR, IT PARQUET CC BY 4.0 Law & Admin. 417.923 561.574 0%
wacky EN, DE, IT TXT CC BY-NC-SA 4.0 Web 8.571.072 6.440 0%
wikihow EN CSV CC BY-NC-SA Knowledge Base 210.526 109.611 0%
wikibooks EU24 XML CC BY-SA 4.0 Books 190.332 168.473 6%
wikinews EU24 XML CC BY 2.5 News 197.265 46.936 3%
wikipedia EU24 XML CC BY-SA 4.0 Knowledge Base 26.175.600 11.382 6%
wikiquote EU24 XML CC BY-SA 4.0 Recreation 208.881 146.497 2%
wikisource EU24 XML CC BY-SA 4.0 Books 428.157 798.322 7%
wikivoyage EU24 XML CC BY-SA 4.0 Culture 71.905 66.278 4%
wikiversity EU24 XML CC BY-SA 4.0 Books 61.397 65.482 0%

### v2

@@ 5.3 Effects of Filtering and Deduplication
Figure 7: Percentage of documents removed after filtering (red) and deduplication (blue) for each language.
In this section, we examine the effects of filtering and deduplication on the dataset at the document level, focusing on their impact across different languages. These analyses provide insights into how preprocessing steps alter the distribution of available data and highlight potential disparities between languages.
Filtering
Filtering involves multiple stages, including prefiltering, quality checks, and harmful perplexity filtering, as illustrated in Figure 2. By comparing the document sizes before and after filtering, we observe that this step significantly reduces the data for all languages, as shown in Figure 7. On average, 96.94% of the documents are removed during filtering.
Notably, filtering exhibits a moderate positive correlation with the size of the original dataset (r=0.530, p=0.0071). This suggests that larger datasets tend to lose a slightly smaller relative percentage of their content during filtering, but the effect is not uniform across languages. Some languages are disproportionately impacted, potentially due to stricter quality control measures or greater prevalence of low-quality or harmful content in smaller datasets.
Deduplication
Our deduplication methodology employs a MinHash + Locality Sensitive Hashing (LSH) approach, as detailed in [64], which optimizes precision and recall in identifying and removing duplicate documents. The deduplication step impacts languages differently, with the percentage of removed documents ranging from 31.50% for English to 67.11% for Bulgarian (Figure 7).
Contrary to intuitive expectations, larger datasets tend to experience a smaller proportion of deduplication. This observation is supported by a strong negative Pearson correlation between the size of the filtered documents and the percentage of data removed during deduplication (r=-0.710, p=0.0001). This suggests that smaller datasets contain a higher proportion of duplicate content relative to their size, which is systematically removed during this stage.
5.3.1 Disparity Index
(a) Deduplication Disparity Index (DDI).
(b) Filtering Disparity Index (FDI).
Figure 8: Disparity Indexes for deduplication and filtering, highlighting languages disproportionately affected by each step.
To quantify the disproportionality in data removal, we introduce the Disparity Index (DI). For each language $l$, the DI is calculated as the Z-score of the removed-to-data ratio $R_{l}$ for each language:
$\operatorname{DI}_{l}=\frac{R_{l}-\mu_{R}}{\sigma_{R}},\quad\text{where}\quad R_{l}=\frac{r_{l}}{D_{l}}$
Here, $r_{l}$ represents the percentage of documents removed for language $l$, $D_{l}$ is the total number of documents for that language, $\mu_{R}$ is the mean of the removed-to-data ratios across all languages, , and $\sigma_{R}$ is the standard deviation. This normalization allows for direct comparisons, with positive and negative Z-scores indicating disproportionate removal or retention, respectively.
Filtering Disparity Index (FDI)
The Filtering Disparity Index ($\operatorname{FDI}$) reveals significant variability in how filtering impacts different languages (Figure 8(b)). Languages with a high positive Z-score, such as Estonian ($\operatorname{FDI}_{et}=2.554$) and Croatian ($\operatorname{FDI}_{hr}=2.186$), lose a disproportionately large portion of their original data during filtering. Conversely, languages with a high negative Z-score, such as Danish ($\operatorname{FDI}_{da}=-1.637$) and English ($\operatorname{FDI}_{en}=-1.141$), are relatively less affected, indicating a smaller proportion of their content is filtered out.
Deduplication Disparity Index (DDI)
Similarly, the Deduplication Disparity Index ($\operatorname{DDI}$) highlights disparities in deduplication across languages (Figure 8(a)). Languages such as Bulgarian ($\operatorname{DDI}_{bg}=1.257$) and Slovak ($\operatorname{DDI}_{sk}=1.174$) are disproportionately impacted, suggesting higher redundancy in their datasets. In contrast, languages such as English ($\operatorname{DDI}_{en}=-3.373$) experience significantly less deduplication, reflecting relatively unique content in their data.
@@ 5.3.2 Summary
Deduplication proved to be the most resource-intensive preprocessing step, consuming 80.8% of the total CPU hours allocated. The introduction of the Deduplication Disparity Index (DDI) and Filtering Disparity Index (FDI) provides a systematic framework for evaluating fairness in the data pipeline. These metrics identify languages disproportionately affected by filtering and deduplication, underscoring the need for tailored processing strategies to ensure equitable data quality and representation across languages.
@@ 8.3.2 List of Curated Datasets
ampsmath_khan [44] EN JSON / TXT MIT Math 102.985 12.081 0%
ampsmath_mathematica [44] EN JSON / TXT MIT Math 4.824.189 232.502 0%
cdrs_bt [21] DE TXT CC0 1.0 Law & Admin. 124.487 515.053 8%
ce_bag [20] DE TXT CC0 1.0 Law & Admin. 5.290 18.487 0%
ce_bfh [29] DE TXT CC0 1.0 Law & Admin. 10.310 22.635 0%
ce_bgh [25] DE TXT CC0 1.0 Law & Admin. 71.976 104.002 0%
ce_bpatg [26] DE TXT CC0 1.0 Law & Admin. 30.772 66.029 0%
ce_bverfg [23] DE TXT CC0 1.0 Law & Admin. 8.449 20.679 0%
ce_bverwg [27] DE TXT CC0 1.0 Law & Admin. 27.099 49.664 0%
cpp_bt [22] DE TXT CC0 1.0 Law & Admin. 4.087 235.502 0%
@@
cd_icj [28] EN, FR TXT CC0 1.0 Law & Admin. 4.565 25.979 0%
cd_pcij [24] EN TXT CC0 1.0 Law & Admin. 518 2.117 0%
dagw [17] DA TXT CC BY 4.0 Web 285.634 922.700 5%
@@
dcep [73] EU24 TXT EUPL Law & Admin. 1.227.996 1.117 5%
medi_notice [36] DE, FR, IT TSV CC BY SA Medical 22.219 52.068 0%
enc2021 [55] EU24 VERT CC BY-NC 4.0 Web 4.361.539 1.004 5%
@@
emea [89] EU24 TXT CC BY 4.0 Medical 17.960 152.252 4%
europarl [54] EU24 XML CC0 1.0 Law & Admin. 63.937 660.716 0%
eurlex [62] EU24 JSONL CC BY 4.0 Law & Admin. 4.463.480 13.235 8%
ecb_corpus [89] EU24 XML CC0 Finance 19 53.894 0%
@@
german_legal_cases [78] DE JSONL CC0 1.0 Law & Admin. 249.240 749.060 0%
german_political_speeches [6] DE XML CC BY-SA 4.0 Law & Admin. 6.659 11.366 0%
greek_legal_code [79] EL JSONL CC BY 4.0 Law & Admin. 40.929 28.791 1%
@@
korpus_malti [71] MA JSONL CC BY-NC-SA Web 103.874 276.204 11% [^29] Noisy data and short document.
@@
marcell [90] BG, HR, HU, RO, SL, SK, PL CONLLUP / TXT CC0 1.0 Law & Admin. 280.629 373.283 5%
macocu [9] EU24 XML CC0 1.0 Web 79.164.225 27.930 12% [^30] Short documents.
multi_un [18] EU24 XML CC BY-NC Law & Admin. 262.391 1.096 0%
multi_eurlex [38] EU24 JSONL CC BY-NC-SA 4.0 Law & Admin. 1.038.305 1.184 1%
medical_t5 [34] DE, EN, ES, FR JSON Apache-2.0 Medical 94.269 90.741 1%
ncc [60] NO JSON Various [^31] The ncc corpus includes the following licenses : NLOD 2.0, CC0 1.0, CC BY-NC 2.0, CC BY-SA 3.0 Knowledge Base 1.241.694 2.920 60% [^32] Short documents.
openlegaldata [78] DE JSONL CC0 1.0 Law & Admin. 103.870 363.962 0%
opensubtitles2018 [66] EU24 XML CC BY-NC Culture 2.257.513 1.398 22% [^33] Noisy documents.
open_web_math [80] EN Parquet ODC-By 1.0 Math 5.956.611 7.378 2%
paisa [68] IT TXT / XML CC BY-NC-SA 3.0 Web 318.730 207.382 7%
parlamint [19] EU24 XML / TEI CC BY Law & Admin. 40.995 1.095 3%
pes2o [87] EN JSONL ODC-By 1.0 Academic 38.958.175 42.172 0%
pile_v2_freelawopinions [32] EN JSONL CC BY-ND 4.0 Law & Admin. 4.410.012 10.408 0%
pile_hackernews [32] EN JSONL CC BY-NC Web 331.205 246.149 1%
pile_nih_exporter [32] EN JSONL CC0 1.0 Medical 930.048 307.055 0%
pile_openwebtext2 [32] EN JSONL MIT Books 15.752.217 10.633 1%
pile_v2_philarchive [32] EN JSONL Various [^34] This Pile part is extracted from ArXiv where the author can decide between a variety of licenses. Academic 40.344 487.032 1%
pile_pmc_abstracts [32] EN JSONL Various [^35] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 15.476.085 3.173 0%
pile_pmc_extracts [32] EN JSONL Various [^36] This Pile part is extracted from PubMed where the author can decide between a variety of licenses. Medical 2.808.849 12.113 1%
ppc [75] PL XML CC BY 4.0 Law & Admin. 38.304 472.193 1%
@@
rp_arxiv [15] EN JSONL Apache-2.0 Academic 1.503.469 10.202 3%
setimes_corpus [88] BG, BS, EL, EN, HR, MK, RO, SQ TXT CC BY-SA 3.0 News 246.206 69.279 7%
seimas [72] EN, LT XML / TMX CC BY 4.0 Law & Admin. 3540 47.319 0%
@@
slwac [67] SL CONLLU CC BY-SA 4.0 Web 1.484.546 953.803 16% [^38] Low language score due to similar languages.
spanish_legal_corpora [40] ES TXT CC BY 4.0 Law & Admin. 15 1.383.749 0%
@@
starcoder_data [65] XX JSONL Various [^40] The Stack is a collection of source code from repositories with various licenses. Any use of all or part of the code gathered in The Stack must abide by the terms of the original licenses, including attribution clauses when relevant. Source Code 206.634.734 73.064 0%
swiss_judgment_prediction [74] DE, FR JSONL CC BY-SA 4.0 Law & Admin. 262.789 144.599 0%
swiss_legislation_corpus [46] DE, FR TXT CC BY-SA Law & Admin. 2 10.407 0%
swiss_policy_documents [36] DE, FR, IT PARQUET CC BY 4.0 Law & Admin. 417.923 561.574 0%
wacky [8] EN, DE, IT TXT CC BY-NC-SA 4.0 Web 8.571.072 6.440 0%
wikihow [56] EN CSV CC BY-NC-SA Knowledge Base 210.526 109.611 0%
wikibooks [30] EU24 XML CC BY-SA 4.0 Books 190.332 168.473 6%
wikinews [30] EU24 XML CC BY 2.5 News 197.265 46.936 3%
wikipedia [30] EU24 XML CC BY-SA 4.0 Knowledge Base 26.175.600 11.382 6%
wikiquote [30] EU24 XML CC BY-SA 4.0 Recreation 208.881 146.497 2%
wikisource [30] EU24 XML CC BY-SA 4.0 Books 428.157 798.322 7%
wikivoyage [30] EU24 XML CC BY-SA 4.0 Culture 71.905 66.278 4%
wikiversity [30] EU24 XML CC BY-SA 4.0 Books 61.397 65.482 0%